OpenStudy (andreaa):

plot the points A= (5,3), B= (-4, -6) and D =(5,-6) on a set of axes. use a ruler to connect them inorder , including D back to A, to form a QUADRILATERAL ( a shape with four sides ) A) what kind of quadrilateral was formed? B) how long is each side of quadrilateral? C) what is the area of the quadrilateral? D) what is the perminitor?

OpenStudy (anonymous):

use the grid paper to draw point D= (-4,3) A) square B) distance use:√(y^2 -y^1)^2 +(x^2-x^1)^2 each side 9 unit C) A = 9(9)=81 D ) P = (9+9)2=36

OpenStudy (andreaa):

idont have any graph paper :(